xx NullConstr

NullConstr

Bela Szilagyi <bela.szilagyi@jpl.nasa.gov>

N T Bishop <bishont@unisa.ac.za> Chi Wai Lai <laicw@astro.unisa.ac.za>

September 09 2005

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11calculatexconstraints 10Scope: restrictedx01  BOOLEAN

11Description: Shall the constraints be calculated?

11                 10  x01  Default: yes

11

11nxangxevxoutsidexeq 10Scope: shared from NULLGRIDx01INT

11

11nxangxghostxpts 10Scope: shared from NULLGRIDx01INT

11

11nxangxptsxinsidexeq10Scope: shared from NULLGRIDx01INT

11

11nxangxstencilxsize 10Scope: shared from NULLGRIDx01INT

11

11nxradialxpts  10Scope: shared from NULLGRIDx01INT

11

7 Interfaces

General Implements: nullconstr

Inherits: nullgrid nullvars nullinterp

Grid Variables 7.0.1 PUBLIC GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 nullxr00 10 NullxR00 x01 compact110
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01  tags11tensortypealias=”Scalar”
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11NullGrid::Nxradialxpts
11                 10  x01variable type11REAL
11 nullxr01 10 NullxR01 x01 compact110
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01  tags11tensortypealias=”Scalar”
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11NullGrid::Nxradialxpts
11                 10  x01variable type11REAL
11 nullxr0a 10 NullxR0A x01 compact110
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01  tags11tensortypealias=”Scalar”
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11NullGrid::Nxradialxpts
11                 10  x01variable type11COMPLEX
11 nullxr0axsplit 10  x01  compact110
11                 10  NullxR0Axr  x01 description11Real and Imaginary part of NullxR0A
11                 10  NullxR0Axi  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01  tags11tensortypealias=”Scalar”
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11NullGrid::Nxradialxpts
11                 10  x01variable type11REAL
11 nullxauxconstr 10 x01 compact110
11                 10NullxAuxConstrxnucnx01 description11Constraints associated with the auxiliary variables
11                 10NullxAuxConstrxcbcnx01 dimensions112
11                 10NullxAuxConstrxckcnx01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01  tags11tensortypealias=”Scalar”
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11NullGrid::Nxradialxpts
11                 10  x01variable type11COMPLEX
11 nullxr00xsouth 10 NullxR00xsouth x01  compact110
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01  tags11tensortypealias=”Scalar”
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11NullGrid::Nxradialxpts
11                 10  x01variable type11REAL

11 x Group Names x 10x Variable Names xx01 Details x11x
11 nullxr01xsouth 10 NullxR01xsouth x01  compact110
11                  10  x01 dimensions112
11                  10  x01 distribution11DEFAULT
11                  10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                  10  x01 group type11ARRAY
11                  10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x01  tags11tensortypealias=”Scalar”
11                  10  x01 timelevels111
11                  10  x01vararrayxsize11NullGrid::Nxradialxpts
11                  10  x01variable type11REAL
11 nullxr0axsouth 10 NullxR0Axsouth x01  compact110
11                  10  x01 dimensions112
11                  10  x01 distribution11DEFAULT
11                  10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                  10  x01 group type11ARRAY
11                  10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x01  tags11tensortypealias=”Scalar”
11                  10  x01 timelevels111
11                  10  x01vararrayxsize11NullGrid::Nxradialxpts
11                  10  x01variable type11COMPLEX
11nullxauxconstrxsouth10 x01 compact110
11                  10NullxAuxConstrxnucsx01 description11Constraints associated with the auxiliary variables (south)
11                  10NullxAuxConstrxcbcsx01 dimensions112
11                  10NullxAuxConstrxckcsx01 distribution11DEFAULT
11                  10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                  10  x01 group type11ARRAY
11                  10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                  10  x01  tags11tensortypealias=”Scalar”
11                  10  x01 timelevels111
11                  10  x01vararrayxsize11NullGrid::Nxradialxpts
11                  10  x01variable type11COMPLEX

11

8 Schedule This section lists all the variables which are assigned storage by thorn PITTNullCode/NullConstr. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxANALYSIS (conditional) x nullconstrxdriver x calculate constraint components r00, r01, r0a x

11x11Language:11fortran
11 x 11Storage: 11nullxr00
11 x 11x  11nullxr01
11 x 11x  11nullxr0a
11 x 11x  11nullxr00xsouth
11 x 11x  11nullxr01xsouth
11 x 11x  11nullxr0axsouth
11 x 11x  11nullxr0axsplit
11 x 11Sync:  11nullxr00
11 x 11x  11nullxr01
11 x 11x  11nullxr0a
11 x 11x  11nullxr00xsouth
11 x 11x  11nullxr01xsouth
11 x 11x  11nullxr0axsouth
11 x 11x  11nullxr0axsplit
11 x 11Triggers: 11nullxr00
11 x 11x  11nullxr01
11 x 11x  11nullxr0a
11 x 11x  11nullxr00xsouth
11 x 11x  11nullxr01xsouth
11 x 11x  11nullxr0axsouth
11 x 11x  11nullxr0axsplit
11 x 11Type: 11function

11

CCTKxANALYSIS (conditional) x nullconstrxauxconstr x calculate constraints associated with auxiliary variables x

11x11Language:11fortran
11 x 11Storage: 11nullxauxconstr
11 x 11x  11nullxauxconstrxsouth
11 x 11Sync:  11nullxauxconstr
11 x 11x  11nullxauxconstrxsouth
11 x 11Triggers: 11nullxauxconstr
11 x 11x  11nullxauxconstrxsouth
11 x 11Type: 11function